OpenStudy (anonymous):

Solve for X. x2– 1 = 0 (1 point) A. x = 0 B. x = –1 and x = 1 C. x = 1 D. x = –1

OpenStudy (angela210793):

x^2-1=(x-1)*(x+1)=0 so x-1=0 or x+1=0 x=1 or x=-1 so is B